OpenStudy (anonymous):

evaluate. integral [ ( 4x^3 ) / sqr(x^2 + 4) ] dx

OpenStudy (kirbykirby):

\[ \int\frac{4x^3}{\sqrt{x^2+4}}dx\] Try this: \[ \int\frac{2x(2x^2)}{\sqrt{x^2+4}}dx\] Let \(x^2+4 = u \implies x^2=u-4\) \(du = 2x\, dx\) So by substitution: \[ \int \frac{2(u-4)}{\sqrt{u}}\,du\]

OpenStudy (anonymous):

sorry, i'm trying to use trig sub

myininaya (myininaya):

If you could do trig sub, but @kirbykirby 's way is actually tons easier. You want to base your trig sub off of that thing in the square root, the x^2+4 thing.

OpenStudy (kirbykirby):

This trick is easier than trig sub I find :) , you can just separate the integral above in each separate integral should be easy to evaluate

OpenStudy (anonymous):

i agree, but the instructions say use trig sub :(

myininaya (myininaya):

It helps me to recall some of the trig identities (all of them is just a rewrite of the some identity, the Pythagorean identitity: 1-sin^2(x)=cos^2(x) 1+tan^2(x)=sec^2(x) sec^2(x)-1=tan^2(x)

myininaya (myininaya):

Do you see what trig identity will be useful here?

myininaya (myininaya):

\[x^2+4=4(\frac{x^2}{4}+1)=4 ( (\frac{x}{2})^2+1)\] And it might help if I factor out that 4 Let me rewrite that thing in the ( ) one more time: \[1+(\frac{x}{2})^2 \] compare this to all of those left sides of the identities I wrote above 1-sin^2(x) 1+tan^2(x) sec^2(x)-1 which one looks the most similar

OpenStudy (anonymous):

x = atan(theta) dx = asec^2(theta)d(theta)

OpenStudy (anonymous):

i have to go, thank you for your help

myininaya (myininaya):

Since 1+tan^2(theta) would be the same as 1+(x/2)^2 if tan(theta)=x/2